# GAST: Sequential Gaussian Avatars with Hierarchical Spatio-temporal Context

### v1

@@ 4.1 Non-Rigid Deformation with Hierarchical Spatio-temporal Context
where $\mathbf{KNN}(\mathbf{T},\mathbf{x}_{i})$ denotes the $\tau$ nearest SMPL template vertexes of the canonical Gaussian position $\mathbf{x}_{i}$. Then we apply an MLP $\mathcal{E}_{\mathcal{V}}$ to encode each Gaussian primitive’s sequential motion state $\mathcal{V}=\{e^{t}|t\in\mathcal{T}\}$ into a point-wise sequential condition $f_{\mathcal{V}}\in\mathbb{R}^{64}$ for non-rigid deformation.
Non-Rigid Deformation. Given the coarse skeleton motion $f_{\Delta\mathcal{P}}=\mathcal{E}_{\Delta\mathcal{P}}(\Delta\mathcal{P})$ condition and fine point-wise velocity condition $f_{\mathcal{V}}=\mathcal{E}_{v}(\mathcal{V})$, we utilize an MLP to predict each Gaussian’s non-rigid deformation as
@@
4.2 Multi-scale Sequence Sampling
To capture both the overall motion trend and the inter-frame motion details, we adopt a multi-scale sequence sampling strategy. Specifically, we sample a series of motion sequences at several progressively increasing intervals to get
$\mathcal{S}=\{s=s_{0}+i\Delta s\}_{i=0}^{i=m},$ (16)
where $\Delta s$ is the increasing rate of sampling interval $s$. We then input the multi-scale sampled sequence motions into the encoder $\mathcal{E}_{\Delta\mathcal{P}}$ and $\mathcal{E}_{\mathcal{V}}$ to obtain hierarchical temporal motion embeddings
$f_{\Delta\mathcal{P}}$ $=\mathcal{E}_{\Delta\mathcal{P}}(\{\Delta\mathcal{P}_{s}\}),$ (17)
$f_{\mathcal{V}}$ $=\mathcal{E}_{\mathcal{V}}(\{\mathcal{V}_{s}\}),$ (18)
where $s\in\mathcal{S}$. In practice, we concatenate all skeleton motion conditions, $\Delta\mathcal{P}_{s}$ for $s\in\mathcal{S}$, and localized vertex motion conditions, $\mathcal{V}_{s}$ for $s\in\mathcal{S}$, across different sampling scales. These are then input to $\mathcal{E}_{\Delta\mathcal{P}}$ and $\mathcal{E}_{\mathcal{V}}$, respectively. More details are shown in the supplementary material.
@@ 4.3 Optimization
where $\mathbf{R}^{\prime}$ is the rotation matrix derived from non-rigid deformed Gaussian’s rotation quaternion $\mathbf{r}^{\prime}$ and $\mathbf{R_{o}}$ represents the rotation matrix in observation space. Following the previous method [17], we utilize an MLP $\mathcal{E}_{lbs}$ to predict the LBS weight offsets for each query canonical point and update the sampled weights from the nearest SMPL vertex as
@@
During the optimization process, we employ four loss functions as supervision, summarized as
@@
where $\lambda$ is used to balance the weight of different losses. $\mathcal{L}_{mask}$ [17] is an $L_{2}$ loss between the rendered $\alpha$ and human body mask. Additional details of the pipeline are provided in the supplementary material.
@@ 5.1 Datasets
DNA-Rendering [6] is a challenging realistic human-centric rendering dataset. It contains diverse scenes from everyday life to professional occasions. We choose 6 typical sequences with loose-fitting garments (1_0206_04, 2_0007_04, 2_0019_10, 2_0044_11, 2_0051_09, 2_0813_05) for experiments. We adopt 24 views for training and 6 views for testing. Novel view rendering results compared with other baselines are reported.
I3D-Human [5] comprises multi-view video frames of humans with loose clothing and complex movements. Compared to previous datasets such as ZJU-MoCap [43], which are mainly collected under controlled speeds and tight-fitting garments, I3D-Human contains scenes closer to daily life. We conduct experiments on 4 sequences following the data split in Dyco [5] and all training and testing data are retained to be the same across all comparison methods. For each sequence, we use 4-5 views for training and the rest for testing.
Table 1: Quantitative Results on DNA-Rendering Dataset. We report the performance of novel view rendering. Our GAST outperforms previous state-of-the-art human modeling methods. We mark the best and the second best methods in cells. LPIPS*=LPIPS ×$\times 10^{3}$.
@@
GauHuman [17] 28.49 0.9534 46.45 28.32 0.9501 52.16 32.55 0.9716 27.91 31.02 0.9670 28.85 26.24 0.9610 38.21 33.79 0.9776 25.14
GAST(Ours) 30.81 0.9649 39.47 29.63 0.9566 43.72 34.97 0.9777 23.53 32.62 0.9747 23.83 28.45 0.9685 34.06 35.81 0.9843 20.83
Figure 3: Qualitative Results on DNA-Rendering Dataset [6]. We compare the visual quality of novel view rendering with previous SOTA methods. To present the differences more clearly, we zoom into the local region and compute the error maps compared with ground truth images. The results show that our method achieves competitive results on both the overall and local region qualities.
Table 2: Quantitative Results on I3D-Human Dataset. We report the performance of novel view rendering. The proposed method achieves competitive results compared with Dyco [5], GauHuman [17], and 3DGS-Avatar [44].
@@
GAST(Ours) 31.81 0.9652 28.64 31.88 0.9636 30.01 31.35 0.9669 34.04 33.47 0.9633 37.90
Figure 4: Qualitative Results on I3D-Human Dataset [5]. We compare the visual quality of novel view rendering with SOTA methods.
@@ 5.2 Baselines and Metrics
We compare against the state-of-the-art human modeling methods including Dyco [5], 3DGS-Avatar [44], GART [27], and GauHuman [17]. Dyco is an implicit NeRF-based method that adopts pose sequence information for temporal modeling. We compare performance with Dyco to showcase the better efficacy of our point-based fine-grained spatio-temporal context design. 3DGS-Avatar, GART, and GauHuman are explicit 3DGS-based methods, and all of them are originally designed for monocular inputs. For fairness, we extend them to multi-view input under the same settings. We compare these methods to demonstrate that finer rendering details can be achieved with our hierarchical sequential conditions for Gaussian points.
We report three key metrics: peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) [48], and learned perceptual image patch similarity (LPIPS) [52].
@@ 5.3 Comparison
Comparison on DNA-Rendering. Tab. 1 shows the quantitative results on DNA-Rendering dataset. The proposed GAST outperforms previous state-of-the-art human modeling methods across all metrics. To demonstrate the visual improvement, we compare the quality of rendering images in Fig. 3. Previous methods, which lack the temporal motion information for non-rigid deformations, are unable to predict the movement of Gaussian primitives properly, resulting in blurred renderings in regions with complex textures (scene 1_0206_04 in Fig. 3). Besides, the results in Fig. 3 show that the compared methods fail to capture the proper shape variation caused by movement (see the white dress in scene 2_0051_09). Although these methods achieve a smooth appearance in these areas, the error maps show that they fail to closely match the actual shape. We leverage the spatio-temporal context to predict Gaussian deformations, allowing our method to capture more accurate detailed appearance variations caused by human motions.
Comparison on I3D-Human. The quantitative results in Tab. 2 show that our method surpasses the previous SOTA method on most metrics, demonstrating our competitive performance on the I3D-Human dataset. Previous NeRF-based Dyco [5] utilizes pose variation as a condition to model human motions. However, it is limited to capturing only overall body motion trends and lacks the capacity for finer-grained modeling in local regions. For example, in ID1_1 and ID2_1 shown in Fig. 4, Dyco fails to model the information in regions far from the human body skeleton, leading to rendering results that do not accurately reflect the true motion. Our method designs more fine-grained conditions for the explicit point-based Gaussian representation to excavate both spatial and temporal motion information hierarchically. The rendering details in Fig. 4 demonstrate the improvement benefiting from such design in these local regions. More results are shown in supplementary materials.
Table 3: Quantitative Ablation Results on I3D-Human Dataset. The Vanilla Non-Rigid MLP in (b) only utilizes Gaussian positions $\mathbf{x}$ and the current frame pose $P$ as the conditions.
@@
(a) Baseline 29.88 0.9570 38.12
(b) (a)+Vanilla Non-Rigid MLP. 31.18 0.9614 35.78
(c) (b)+$\Delta\mathcal{P}$ Conditions. 31.89 0.9640 33.86
(d) (c)+ $\mathcal{V}$ Conditions. 32.05 0.9646 32.82
(e) (d)+Multi-scale Sampling. (Ours) 32.13 0.9647 32.65
@@ 5.4 Ablation Study
Multi-scale sequence sampling is designed to capture both the overall motion trend and the fine-grained inter-frame motion details. With temporal conditions sampled at different intervals, our method is able to predict each Gaussian’s deformation more robustly, leading to better rendering quality as shown in Tab. 3 (e) and Fig. 5 (e).
@@ 6 Limitations and Conclusion
Limitations. Compared to prior NeRF-based approaches, our method models the human body with the point-based Gaussian representation, which may introduce needle-like artifacts during rendering. As indicated in Tab. 2, the metric on LPIPS is slightly inferior to that of NeRF-based Dyco [5]. We notice that some subsequent works [19, 15] on 3DGS optimize the artifacts by introducing surface-related smooth regularization. Generative model-based human avatar [26] could also potentially be introduced as a guarantee of high-fidelity rendering. However, these are beyond the scope of our motion modeling and will serve as future work to improve rendering quality.
Conclusion. In this paper, we propose GAST, a 3DGS-based framework that integrates hierarchical spatial-temporal information for 3D human modeling. Specifically, we introduce a coarse-to-fine motion context for the non-rigid deformation of Gaussian primitives. At the spatial dimension, the global and local region transformations for Gaussians are learned based on the overall inter-frame human pose variations and point-wise motions. To capture stable temporal motion information for each point, we introduce an SMPL vertex velocity template and sample the velocity of each Gaussian point from it. Along the temporal dimension, we apply a multi-scale sampling strategy to capture both long-term motion trends and inter-frame details, supporting more robust predictions of Gaussian deformations. Through the above design, we have achieved state-of-the-art rendering of human avatars with complex garments and motions.
@@ 7 Details of Skeleton Motion $\Delta P$
The pose $P$ is represented as the rotations of all joints in axis-angle form:
@@ 8 Datasets
I3D-Human Dataset [5] We use images with resolution 512 $\times$ 512 for experiments following Dyco [5].
DNA-Rendering Dataset [6] We use the original images’ resolution 512 $\times$ 612 for experiments. For the dataset split, we sample 24 evenly distributed views from camera IDs 0–48 for training and 6 additional uniformly distributed views from IDs 48–60 for novel view testing.
@@ 9 Motion Embeddings for Non-Rigid MLP
The embeddings $f_{\Delta\mathcal{P}}$ and $f_{\mathcal{V}}$ in Eq. (12) describe the skeleton and local region motions, respectively. For each frame, all Gaussian primitives $\{\mathcal{G}_{i}\}$ share the same skeleton motion embedding $f_{\Delta\mathcal{P}}\in\mathbb{R}^{32}$ and we concat it with $\mathcal{G}_{i}$’s point-wise motion embedding ${f_{\mathcal{V}}}_{i}\in\mathbb{R}^{64}$ as a new condition $f_{i}^{\prime}\in\mathbb{R}^{96}$. $f_{i}^{\prime}$ is then input to a 2-layer MLP $\mathcal{E}_{non-rigid}$ to predict the corresponding non-rigid deformation.
@@ 10 Details of Loss Functions
We use $\mathcal{L}_{ssim}$ to constrain the structure similarity between rendered image and the ground truth, which is given by
@@
where $\text{SSIM}(\cdot)$ is the SSIM metric. Additionally, we use the LPIPS loss to ensure the perceptual similarity:
@@
where $\text{LPIPS}(\cdot)$ is the LPIPS metric. Following GauHuman [17] and 3DGS-Avatar [44], we employ a mask loss to ensure that Gaussian primitives are accurately localized within their designated regions:
@@
Figure 6: Additional Novel View Qualitative Results on DNA-Rendering Dataset [6]. We provide both the complete image and localized error map comparisons against other methods for a more comprehensive visualization.

### v2

@@ 4.2 Spatio-temporal Multi-scale Sampling (STMS)
where $\mathbf{KNN}(\mathbf{T},\mathbf{x}_{i})$ denotes the $\tau$ nearest SMPL template vertexes of the canonical Gaussian position $\mathbf{x}_{i}$. Similar to Eq. 7, we then apply an MLP $\mathcal{E}_{\mathcal{V}}$ to encode each Gaussian primitive’s sequential motion embedding $\mathcal{V}=\{e^{t}|t\in\mathcal{T}\}$ into a point-wise sequential condition $f_{\mathcal{V}}\in\mathbb{R}^{96}$
$f_{\mathcal{V}}=\mathcal{E}_{\mathcal{V}}(\mathcal{V}).$ (13)
In the temporal dimension, to capture both the overall motion trend and the inter-frame motion details, we adopt a multi-scale sequence sampling strategy. In detail, we sample a series of sequences at several progressively increasing intervals to get human motion information across different temporal windows
$\mathcal{S}=\{s=s_{0}+i\Delta s\}_{i=0}^{i=m},$ (14)
where $\Delta s$ is the increasing rate of sampling interval $s$. We then input the multi-scale sampled sequence motions into the skeleton motion encoder $\mathcal{E}_{\Delta\mathcal{P}}$ and the point-wise motion encoder $\mathcal{E}_{\mathcal{V}}$ to obtain hierarchical temporal motion embeddings. Therefore, Eq. 7 and Eq. 13 can be revised as
$f_{\Delta\mathcal{P}}$ $=\mathcal{E}_{\Delta\mathcal{P}}(\{\Delta\mathcal{P}_{s}\}),\quad s\in\mathcal{S}$ (15)
$f_{\mathcal{V}}$ $=\mathcal{E}_{\mathcal{V}}(\{\mathcal{V}_{s}\}),\quad s\in\mathcal{S}.$ (16)
In practice, we concatenate all skeleton motion conditions, $\Delta\mathcal{P}_{s}$ for $s\in\mathcal{S}$, and localized vertex motion conditions, $\mathcal{V}_{s}$ for $s\in\mathcal{S}$, across different sampling scales. These are then input to $\mathcal{E}_{\Delta\mathcal{P}}$ and $\mathcal{E}_{\mathcal{V}}$, respectively. More details are shown in the supplementary material.
Non-Rigid Deformation. Given the coarse skeleton motion $f_{\Delta\mathcal{P}}=\mathcal{E}_{\Delta\mathcal{P}}(\{\Delta\mathcal{P}_{s}\})$ condition and fine point-wise velocity condition $f_{\mathcal{V}}=\mathcal{E}_{\mathcal{V}}(\{\mathcal{V}_{s}\})$, we utilize an MLP to predict each Gaussian’s non-rigid deformation as
@@ 4.3 Optimization
where $\mathbf{R}^{\prime}$ is the rotation matrix derived from the non-rigid deformed Gaussian’s rotation quaternion $\mathbf{r}^{\prime}$, and $\mathbf{R_{o}}$ represents the rotation matrix in observation space. Following the previous method [22], we utilize an MLP $\mathcal{E}_{lbs}$ to predict the LBS weight offsets for each query canonical point and update the sampled weights from the nearest SMPL vertex as
@@
During the optimization process, we employ a combination of loss functions as supervision, summarized as
@@
where $\lambda$ is used to balance the weight of different losses. $\mathcal{L}_{mask}$ [22] is an $L_{2}$ loss between the rendered $\alpha$ and human body mask. Similar to [56], we apply $\mathcal{L}_{isopos}$ and $\mathcal{L}_{isocov}$ to control the Gaussian primitive’s position and covariance. Details are shown in the supplementary material.
@@ 5.1 Datasets
DNA-Rendering [9] is a challenging human-centric rendering dataset. It contains diverse scenes from everyday life to professional occasions. We use 6 sequences with loose-fitting garments (1_0206_04, 2_0007_04, 2_0019_10, 2_0044_11, 2_0051_09, 2_0813_05) for experiments. We adopt 24 views for training and 6 views for testing.
I3D-Human [8] contains scenes closer to daily life, which comprises multi-view video frames of humans with loose clothing and complex movements. We conduct experiments on 4 sequences and use 4-5 views for training and the rest for testing. All training and testing data are retained to be the same in comparisons following Dyco’s [8] data split.
Table 1: Quantitative Results on DNA-Rendering Dataset. We report the performance of novel view rendering. Our method outperforms previous state-of-the-art human modeling methods. We mark the best and the second best methods in cells. LPIPS*=LPIPS ×$\times 10^{3}$.
@@
GauHuman [22] 27.96 0.9500 50.42 27.93 0.9496 56.93 31.63 0.9698 32.82 30.38 0.9641 34.19 25.99 0.9599 42.00 33.43 0.9764 29.42
Ours 30.81 0.9649 39.47 29.63 0.9566 43.72 34.97 0.9777 23.53 32.62 0.9747 23.83 28.45 0.9685 34.06 35.81 0.9843 20.83
Figure 3: Novel View Qualitative Results on DNA-Rendering. We zoom into the local region and compute the error maps compared with ground truth images. The results show that our method achieves competitive results on both the overall and local region qualities.
Table 2: Novel View Rendering Quantitative Results on I3D-Human.
@@
Ours 31.81 0.9659 27.03 31.98 0.9660 27.84 31.53 0.9685 30.23 33.62 0.9651 34.02
Figure 4: Novel Pose Qualitative Results on I3D-Human. We compare the proposed method with previous SOTA approaches. The rendered images and error maps demonstrate our robustness on novel pose rendering.
@@ 5.2 Baselines and Metrics
We compare against the state-of-the-art human modeling methods including Dyco [8], 3DGS-Avatar [56], GART [31], and GauHuman [22]. Dyco is an implicit NeRF-based method that adopts pose sequence information for temporal modeling. We compare performance with Dyco to showcase the better efficacy of our fine-grained motion condition design. 3DGS-Avatar, GART, and GauHuman are explicit 3DGS-based methods, and all of them are originally designed for monocular inputs. For fairness, we extend them to multi-view input under the same settings. We report three key metrics: peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) [66], and learned perceptual image patch similarity (LPIPS) [77].
@@ 5.3 Comparison
Figure 6: Out-of-distribution Pose Animation.
Comparison on DNA-Rendering. Tab. 1 shows the quantitative results on DNA-Rendering dataset. The proposed method outperforms previous state-of-the-art human modeling methods across all metrics. To demonstrate the visual improvement, we compare the quality of rendering images in Fig. 3. Previous methods, which lack the temporal motion information for non-rigid deformations, are unable to predict the movement of Gaussian primitives properly, resulting in blurred renderings in regions with complex textures (scene 1_0206_04 in Fig. 3). Besides, the results in Fig. 3 show that the compared methods fail to capture the proper shape variation caused by movement (e.g., the white dress in scene 2_0051_09). Although these methods achieve a smooth appearance in these areas, the error maps show that they fail to match the actual shape closely. We leverage the hierarchical motion context to predict Gaussian deformations, allowing our method to capture detailed appearance variations caused by human motions more accurately.
Comparison on I3D-Human. The quantitative results in Tab. 2 show that our method surpasses the previous SOTA method on most metrics, demonstrating our competitive performance on the I3D-Human dataset. Previous NeRF-based Dyco [8] utilizes pose variation as a condition to model human motions. However, it is limited to capturing only overall body motions and lacks the capacity for finer-grained modeling in local regions. For example, in ID1_1 and ID2_1 shown in Fig. 8 (in the Supp. Mat.), our method renders results that accurately reflect the true motion, while Dyco fails to model the information in regions far from the human body skeleton. Tab. 3 reports the quantitative results of novel pose rendering on the I3D-Human dataset. Our method achieves best performance among 3DGS-based human modeling methods and also enables real-time rendering ($\sim$ 45 FPS on I3D-Human) compared to NeRF-based Dyco ($\sim$ 0.7 FPS). Fig. 4 shows the qualitative comparisons of novel pose rendering. Different from Dyco [8], which relies solely on joint motions as conditions, our method leverages the motion state of each Gaussian primitive to predict deformations, enabling more flexible modeling of motions in regions distant from the human body.
To evaluate the generalizability, we also conduct experiments on ZJU-MoCap [53]. Please refer to the supplementary materials for more details.
Out-of-Distribution Poses. To evaluate the generalization ability to novel poses, we conduct experiments using animations with large pose variations. Specifically, we use a model trained on one sequence to render target poses sampled from other unseen sequences, as different sequences in DNA-Rendering and I3D-Human contain distinct motion patterns. As illustrated in Fig. 6, the Target Pose is drawn from an unseen sequence exhibiting different human motions. The results suggest that our method is capable of generalizing to these significant pose changes, even though such poses were not observed during training.
Table 3: Metrics of Novel Pose Renderings on I3D-Human.
@@
Dyco [8] 30.09 0.9569 35.32
3DGS-Avatar [56] 29.54 0.9555 38.59
GauHuman [22] 29.31 0.9526 48.29
Ours 30.28 0.9583 36.07
@@ 5.4 Ablation Study
Spatio-temporal Multi-scale Sampling (STMS) is designed to enhance the robustness of human motion conditions by incorporating local region motion information and motion patterns across different temporal windows. With such a hierarchical motion context embedding design, our method is able to predict each Gaussian’s deformation more robustly, leading to better rendering quality as shown in Tab. 4 (e) and Fig. 5 (e). Please refer to the supplementary materials for more detailed ablation experiments.
Table 4: Quantitative Ablation Results on I3D-Human. (a) denotes experiments without non-rigid deformation. (b) only utilizes Gaussian positions $\mathbf{x}$ and the current frame pose $P$ as the conditions for non-rigid MLP.
Methods PSNR $\uparrow$ SSIM $\uparrow$ LPIPS* $\downarrow$
(a) Baseline 29.76 0.9569 38.35
[…]
(d) (c)+ $\mathcal{V}$ Conditions. 32.01 0.9651 31.23
(e) (d)+STMS. (Ours) 32.24 0.9664 29.78
@@ 6 Limitations and Conclusion
Limitations. The Gaussian-based representation used in our method may introduce slight blur artifacts during rendering, whereas NeRF’s ray-based volumetric integration tends to produce sharper results. Moreover, our local velocity cues are derived from the coarse SMPL model rather than dense surface tracking [79], which may limit the accuracy of fine-scale garment deformation. Addressing these limitations remains an open challenge for future work.
Conclusion. In this paper, we propose a 3DGS-based framework that integrates hierarchical motion context for 3D human modeling. Specifically, the non-rigid deformation for the Gaussian primitive is learned based on the global human skeleton variations and fine-grained Gaussian’s point-wise motions. To capture each Gaussian primitive’s motion information more robustly, we introduce a spatio-temporal multi-scale sampling strategy to incorporate both local region motion features and motion patterns across different temporal intervals. Through the above design, the proposed method achieves state-of-the-art rendering of human avatars with complex garments and motions.
Acknowledgment
This research was supported by Shanghai Artificial Intelligence Laboratory.
@@ 7 Details of Skeleton Motion $\Delta P$
The pose $P$ is represented as the rotation relative to the parent node of all joints:
@@ 8 Datasets
I3D-Human Dataset [8]. We use images at a resolution of 512 $\times$512 for experiments and follow Dyco[8] for both training and testing splits.
DNA-Rendering Dataset [9]. We use images with a resolution of 512 $\times$ 612 for training and 750 $\times$ 1024 for testing in our experiments. For the temporal split, we select 100 consecutive time steps (pose IDs 0–99) and use the multi-view images of all selected steps for training. During testing, we evaluate on a total of 20 time steps sampled every 5 steps from this training range, using multi-view images from held-out camera views. For the view split, we uniformly sample 24 views from camera IDs 0–48 for training and 6 additional views uniformly sampled from camera IDs 48–60 for novel view evaluation.
ZJU-MoCap [53]. We use images with resolution 512 $\times$ 512 for experiments and follow 3DGS-Avatar[56] for both training and testing splits.
@@ 9 Motion Embeddings for Non-Rigid MLP
The embeddings $f_{\Delta\mathcal{P}}$ and $f_{\mathcal{V}}$ in Eq. (17) describe the skeleton and local region motions, respectively.
For each frame, all Gaussian primitives $\{\mathcal{G}_{i}\}$ share the same skeleton motion embedding $f_{\Delta\mathcal{P}}\in\mathbb{R}^{32}$ and we concat it with $\mathcal{G}_{i}$’s point-wise motion embedding ${f_{\mathcal{V}}}_{i}\in\mathbb{R}^{96}$ as a new condition $f_{i}^{\prime}\in\mathbb{R}^{128}$. $f_{i}^{\prime}$ is then input to an MLP $\mathcal{E}_{non-rigid}$ to predict the corresponding non-rigid deformation. $\mathcal{E}_{non-rigid}$ consists of 3 hidden layers and one output layer, and each hidden layer is followed by a ReLU activation.
@@ 10 Details of Loss Functions
We use $\mathcal{L}_{ssim}$ to constrain the structure similarity between the rendered image and the ground truth, which is given by
@@
[…]
@@
[…]
@@
Similar to [56, 54], we also utilize the as-isometric-as-possible constraint [28] to enforce neighboring distance similarity of 3D Gaussian centers and covariance matrices between canonical space and observation space:
$\mathcal{L}_{isopos}$ $=\sum_{i=1}^{N}\sum_{j\in\mathbf{KNN}(i)}|d(\mathbf{x}^{i},\mathbf{x}^{j})-d(\mathbf{x_{o}}^{i},\mathbf{x_{o}}^{j})|,$ (35)
$\mathcal{L}_{isocov}$ $=\sum_{i=1}^{N}\sum_{j\in\mathbf{KNN}(i)}|d(\mathbf{\Sigma}^{i},\mathbf{\Sigma}^{j})-d(\mathbf{\Sigma_{o}}^{i},\mathbf{\Sigma_{o}}^{j})|,$ (36)
where $N$ is the number of Gaussian primitives. $\mathbf{x},\mathbf{\Sigma}$ and $\mathbf{x_{o}},\mathbf{\Sigma_{o}}$ are the Gaussian primitive’s position and covariance matrix in canonical space and observation space, respectively. $\mathbf{KNN}$ denotes the $K=5$ nearest points.
Figure 7: Novel View Qualitative Results on DNA-Rendering. We provide both the complete image and localized error map comparisons against other methods for a more comprehensive visualization.